Dunder-Mifflin, Inc.​ (DMI) is selling​ 600,000 bonds to raise money for the publication of new magazines in the coming year. The bonds will pay a coupon rate of 14.2​% with semiannual payments and will mature in 30 years. Its par value is ​$100. DMI hires an investment banker for the sale of the​ 600,000 bonds. The investment banker charges a fee of 2​% on each bond sold. What is the cost of debt to DMI if the following are the proceeds before the​ banker's fees are​ deducted? a.  ​$49584000 b.  ​$52890000 c.  ​$64386000 d.  ​$73398000
Answer #1

1
=RATE(30*2,14.2%*100/2,-49584000/(600000)*(1-2%),100)*2=17.5602410870702%

2
=RATE(30*2,14.2%*100/2,-52890000/(600000)*(1-2%),100)*2=16.4604004767674%

3
=RATE(30*2,14.2%*100/2,-64386000/(600000)*(1-2%),100)*2=13.4892846466501%

4
=RATE(30*2,14.2%*100/2,-73398000/(600000)*(1-2%),100)*2=11.7797023157377%